OpenStudy (anonymous):

HELP ME. Which answer is correct? What is the sum of the first five terms of the geometric sequence in which a1 = 5 and r = 1/5? Express your answer as an improper fraction.

OpenStudy (anonymous):

I got 2999/625 Is that right?

OpenStudy (anonymous):

s5=5(1-1/5^5) / 1- 1/5 s5=5(1-0.00032) / 4/5 s5=5(0.99968) / 4/5 s5=4.9984/ 4/5 s5=6.248 781/125 I got a different answer this time.